### v1

# The Susceptibility Paradox in Online Social Influence

Abstract
Understanding susceptibility to online influence is crucial for mitigating the spread of misinformation and protecting vulnerable audiences. This paper investigates susceptibility to influence within social networks, focusing on the differential effects of influence-driven versus spontaneous behaviors on user content adoption. Our analysis reveals that influence-driven adoption exhibits high homophily, indicating that individuals prone to influence often connect with similarly susceptible peers, thereby reinforcing peer influence dynamics. Conversely, spontaneous adoption shows significant but lower homophily. Additionally, we extend the Generalized Friendship Paradox to influence-driven behaviors, demonstrating that users’ friends are generally more susceptible to influence than the users themselves, de facto establishing the notion of Susceptibility Paradox in online social influence. This pattern does not hold for spontaneous behaviors, where friends exhibit fewer spontaneous adoptions. We find that susceptibility to influence can be accurately predicted using friends’ susceptibility alone, while predicting spontaneous adoption requires additional features, such as user metadata. These findings highlight the complex interplay between user engagement and preferences in spontaneous content adoption. Our results provide new insights into social influence mechanisms and offer implications for designing more effective moderation strategies to protect vulnerable audiences.
Introduction
On social media, social influence plays a significant role in disseminating information and shaping beliefs, attitudes, and behaviors. Politically, platforms like Twitter and Facebook are arenas for spreading biased information and fake news, which can potentially disrupt the integrity of voting events (González-Bailón et al. 2023; Stella, Ferrara, and De Domenico 2018). Economically, brands can sway consumers’ purchasing decisions through targeted advertising and influencer partnerships (Aral and Walker 2012; Bakshy et al. 2012a). Socio-culturally, social media is pivotal in accelerating social movements and activism (Den Hond and De Bakker 2007; Mundt, Ross, and Burnett 2018).
Research has shown that online social influence can be a powerful tool for persuading behavioral change, especially when messages are tailored to specific audiences (Oyibo and Vassileva 2019). Identifying groups highly susceptible to such influences is, therefore, essential. While many studies have examined various factors that contribute to susceptibility to influence—such as personality traits (Oyibo and Vassileva 2019), political leanings (Calvillo, Rutchick, and Garcia 2021), information credibility (Traberg and van der Linden 2022), and exposure frequency (Hassan and Barber 2021)—they often overlook the relational aspect and networked nature of influence dynamics. Consequently, the interconnectedness and potential peer influence within networks are not fully accounted for understanding and predicting susceptibility. For instance, how an individual’s likelihood of being influenced relates to the susceptibility levels of their social connections remains underexplored.
Contributions of this work
In this paper, we present an empirical framework designed to understand, model, and predict users’ susceptibility to influence on social media platforms. This framework takes into account the networked nature of social influence, with a focus on homophilous influence dynamics, and investigates the Generalized Friendship Paradox, examining how it manifests in influence-driven versus spontaneous behaviors. In detail, we address the following Research Questions (RQs):
  1. RQ1:
    Is susceptibility to influence homophilous within a social network? We aim to investigate whether a user’s susceptibility is correlated with the susceptibility of their friends.
  2. RQ2:
    Does the Generalized Friendship Paradox extend to user susceptibility to influence? Our objective is to explore whether a user’s friends are typically more susceptible to influence than the user themselves.
  3. RQ3:
    How effectively can a user’s susceptibility be predicted by their friends’ susceptibility? We seek to measure to what extent a user’s level of susceptibility can be predicted using information from their social network.
By leveraging two large-scale Twitter datasets covering political and public health discussions, our study reveals that susceptibility to influence is homophilous within social networks. We demonstrate that influence-driven behaviors exhibit greater homophily than spontaneous behaviors. Additionally, we found that users’ friends are generally more susceptible to influence and less prone to spontaneous actions than the users themselves, supporting that the Generalized Friendship Paradox holds for influence-driven behavior but not for spontaneous action in friendship networks. These insights inform predictive models, which can predict users’ rate of influence-driven adoptions by leveraging only their friends’ susceptibility to influence. In contrast, spontaneous behaviors are more difficult to predict using solely friends’ influence signals, suggesting a complex interplay of user engagement and preferences in spontaneous actions.
Our findings are consistent across the two large-scale Twitter datasets and provide crucial insights into diverse behaviors, vulnerabilities, and influence dynamics on social media platforms. This knowledge can inform network interventions designed to protect vulnerable audiences and mitigate risks associated with influence operations, propaganda, and misinformation campaigns.
Related Work
Susceptibility to Social Influence
Numerous studies have investigated the factors that make individuals susceptible to influence, categorizing these factors as either personal or external. From a personal standpoint, demographics such as age and gender play significant roles, with younger individuals and men often being more susceptible to influence when adopting new products (Aral and Walker 2012). Personality traits like neuroticism, openness, and conscientiousness are also recognized as consistent predictors of susceptibility to influence strategies (Oyibo and Vassileva 2019). In the political sphere, individuals with right-leaning ideologies are more susceptible to misinformation and deception (Calvillo, Rutchick, and Garcia 2021; Luceri et al. 2019; Pennycook and Rand 2019), particularly those who exhibit lower trust in scientific authority or a tendency towards conspiracy thinking (Saling et al. 2021).
External factors also play a role in susceptibility. Individuals are more likely to accept claims from sources they perceive as credible (Traberg and van der Linden 2022), and this tendency is further amplified when the persuasive message aligns with their partisanship or ideological views (Guess, Nyhan, and Reifler 2018; Moravec, Minas, and Dennis 2018). Additionally, the “illusory truth” effect occurs when repeated exposure to specific information makes a claim appear more believable, regardless of its veracity, thereby increasing its acceptance (Hassan and Barber 2021).
Influence and Network Dynamics.
A user network may be an under-explored external factor that affects an individual’s susceptibility to social influence. Individuals with similar characteristics often cluster within networks, leading to homophilous patterns of content exposure and adoption behaviors (McPherson, Smith-Lovin, and Cook 2001; Aral and Walker 2012). Furthermore, the structure of these networks often enhances the visibility of information among similar users, potentially fostering echo chambers (Cinelli et al. 2021). These effects are intensified by algorithmic recommendations that tailor content exposure to match the prevailing interests within one’s network (Guess, Nyhan, and Reifler 2020). Such biased exposure can contribute to the “majority illusion,” where individuals mistakenly believe that the views or behaviors prevalent in their immediate social circle are representative of the general population (Lerman, Yan, and Wu 2016). A structural cause of this phenomenon is known as the “friendship paradox,” which we will explore in the following section.
The Generalized Friendship Paradox
Network structure can skew local observations. The friendship paradox states that, on average, your friends have more friends than you do (Feld 1991). This occurs because high-degree nodes (individuals with many friends) are more likely to be sampled when considering the friends of a random individual. The friendship paradox is more pronounced in networks with heavy-tailed degree distributions, where a small number of nodes have a large number of connections (Lerman, Yan, and Wu 2016).
The Generalized Friendship Paradox (GFP) extends the original concept to node attributes beyond the number of friends. It posits that the characteristics of the neighbors of a node in any network will, on average, be greater than those of the node itself (Eom and Jo 2014). For example, Hodas, Kooti, and Lerman (2013) examined the GFP in the context of online social networks, showing that users’ friends are generally more active and receive more viral content on platforms like Twitter. Additionally, Higham (2019) found that individuals’ friends are more central in social networks.
Research also suggests that the GFP could bias local perceptions and individuals’ understanding of societal norms (Alipourfard et al. 2020). Christakis and Fowler (2013) highlight that the spread of behaviors such as smoking, drinking, and exercising is influenced by perceptions of peers’ behaviors. Similarly, susceptibility to adopting information might increase when individuals perceive them as common among their friends, thereby reinforcing the influence of networks on information-sharing behaviors.
Methodology
Framework for Susceptibility Modeling
Susceptibility to online influence is commonly modeled in probabilistic terms, indicating the likelihood that a user’s behavior and/or beliefs may be influenced by online social interactions (Zhou and Zafarani 2019; Romero, Meeder, and Kleinberg 2011; Goyal, Bonchi, and Lakshmanan 2010). Existing literature consistently highlights two primary components of susceptibility to influence: Exposure to content and the potential subsequent Adoption of that content (Aral and Walker 2012; Hoang and Lim 2016). We adhere to this formulation to construct our Susceptibility Framework, which aims to estimate users’ susceptibility to influence based on the dynamics between exposure and adoption.
Exposure.
On social media, exposure to a piece of content typically refers to a user encountering or viewing the content. Measuring exposure from observational social media data can be challenging, as detailed user behavioral data, such as scrolling and clicking, remain largely inaccessible to academics. Additionally, we don’t have access to the structure of the user follower network and, therefore, whose content each user is exposed to from those whom they follow. The situation is further complicated by algorithmic feed curation, i.e., recommendation algorithms displaying potentially relevant content to a user that does not originate from the users they are following (Guess, Nyhan, and Reifler 2020). Despite these challenges, several studies have approximated exposure by analyzing user interactions (Ferrara and Yang 2015; Rao, Morstatter, and Lerman 2022; Sasahara et al. 2021; Ye et al. 2024). The core assumption is that any form of engagement—whether retweeting, quoting, or replying—indicates a user’s potential exposure to another user’s content. These interactions imply that the user has not only viewed but also interacted with the content, thereby increasing the likelihood of subsequent exposures to that user’s content. The rationale is to leverage past interactions as a proxy for predicting future exposure.
Therefore, we operationalize exposure as follows: If a user $u_{t}$ interacts with another user $u_{s}$ through a retweet, quote, or reply, we consider $u_{t}$ exposed to all tweets posted by $u_{s}$ after their first observed interaction.
Adoption.
In this study, we define adoption as the action of sharing a URL, whether through an original tweet, a retweet, a quoted tweet, or a reply. We focus on URLs because they are commonly embedded in tweets and correspond to specific pieces of information, allowing for straightforward tracking of their diffusion and adoption within a social network. Sharing a URL, whether through original content or interactions, represents a form of information adoption, reflecting a user’s active decision to disseminate specific content. However, it is crucial to note that sharing or adopting a URL does not necessarily imply trust or endorsement of its content. While retweeting is often viewed as an endorsement (Boyd, Golder, and Lotan 2010; Metaxas et al. 2015), other sharing activities can serve to question or critique the content (Hemsley et al. 2018).
Figure 1: An illustration of our susceptibility framework.
Susceptibility Metrics.
Building on our conceptualization of Exposure and Adoption, we develop a framework to estimate user susceptibility to influence. Specifically, by analyzing user activity, we evaluate two crucial dimensions: potential exposure to the content and potential adoption of the content. It is important to acknowledge that our data is ultimately an incomplete rendition of the real world: There may be instances where users have shared or been exposed to content outside of our dataset.
Henceforth, we classify the relationship between a user and a piece of content into four distinct scenarios:
  1. (i)
    Exposed and adopted: the user viewed and then shared the content;
  2. (ii)
    Exposed but not adopted: the user viewed the content but did not share it;
  3. (iii)
    Not exposed but adopted: the user shared the content without any prior exposure;
  4. (iv)
    Neither exposed nor adopted: the user neither viewed nor ever shared the content.
Based on this framework, and in line with findings by Aral and Walker (2012), we categorize adoption into two main types: influence-driven adoption, where the content is shared post-exposure, and spontaneous adoption, where the content is shared without any prior exposure. Recognizing these distinctions is crucial, particularly since adoptions can occur independently of direct influence from others. Motivated by this categorization, we introduce two susceptibility metrics: the Influence-Driven Adoption Rate (IAR) and the Spontaneous Adoption Rate (SAR).
Influence-driven Adoption Rate (IAR). This metric calculates the ratio of URLs that a user adopted after being exposed to them to the total number of URLs the user was exposed to, yielding a value between 00 and $1$. Mathematically, it is expressed as:
$\textbf{IAR}=p(\text{adopted}\mid\text{exposed})=\frac{|E_{u_{t}}\cap A_{u_{t}}|}{|E_{u_{t}}|},$
where $E_{u_{t}}$ denotes the set of all tweets containing URLs to which the user $u_{t}$ was exposed, and $A_{u_{t}}$ denotes the set of all tweets containing URLs that the user $u_{t}$ actually adopted. Following the methodology in (Bakshy et al. 2012b; Ye et al. 2024), we implement a buffer period of two months for $A_{u_{t}}$. This period counts adoptions only if they occur after the first two months of our dataset, ensuring there is enough time to observe potential exposures before any subsequent adoptions.
Spontaneous Adoption Rate (SAR). This metric measures the proportion of URLs that a user adopts without prior exposure, with values ranging from 00 to $1$. It is mathematically expressed as:
$\textbf{SAR}=1-p(\text{exposed}\mid\text{adopted})=1-\frac{|E_{u_{t}}\cap A_{u_{t}}|}{|A_{u_{t}}|},$
where $E_{u_{t}}$ and $A_{u_{t}}$ are defined similarly to the IAR metric.
The distributions of the IAR and SAR metrics for both datasets are shown in Figure 6 (see Appendix). The strong negative Spearman correlation between IAR and SAR (Election dataset: $\rho=-0.855$, $\textit{p}<0.001$; Covid dataset: $\rho=-0.857$, $\textit{p}<0.001$) indicates that users who frequently adopt content after being exposed to it (i.e., high IAR) tend to adopt content without prior exposure less frequently (i.e., low SAR), and vice versa. This observation supports the validity of our susceptibility framework and particularly these two metrics, in capturing distinct behaviors on social media—whether users are more influenced by external sources (high IAR) or more prone to independent (i.e., spontaneous) content sharing (high SAR). It is important to note that a high susceptibility level is typically associated with a high IAR and a low SAR. We maintain these two distinct metrics to more accurately reflect the nuances of influence dynamics and user behaviors.
Data Collection
To ensure the robustness of our findings across different contexts, we conduct our experiments on two datasets from Twitter (now X), referred to as the Election dataset and the Covid dataset. These datasets focus on two major events: the 2020 US Presidential Election and the COVID-19 global pandemic, covering a comprehensive range of topics related to political and public health discussions.
The Election dataset (Chen, Deb, and Ferrara 2021) was collected beginning January 1st, 2020, and the Covid dataset (Chen, Lerman, and Ferrara 2020) beginning January 21st, 2020. For this research, we utilize data from the 12-month period of the same year, ending on December 31, 2020. Both datasets were collected via Twitter’s streaming API, which gathered a sample of all tweets using a list of manually curated, topic-specific keywords, accounts, and hashtags. The tweets were fetched in real-time, ensuring data completeness and avoiding well-known re-hydration issues (Pfeffer et al. 2023).
Table 1: Number of users and tweets, along with the number of nodes ($|V|$) and edges ($|E|$) in the user networks $G(V,E)$.
Election Covid
All tweets w/ URL 64M 76M
Total users 4.5M 10.8M
Target users 0.7M 1.0M
Original tweets 7.2M 18.8M
Retweets 27.5M 38.4M
Quoted tweets 18.5M 10.9M
Reply tweets 10.8M 7.9M
$G(V,E)$ $|V|$ $|E|$ $|V|$ $|E|$
Interaction 49,794 117,384 73,960 163,974
Retweet 25,077 49,679 53,161 90,800
Mention 45,289 72,554 56,060 81,969
Since our methodology relies on tracking users’ exposure to and adoption of information via URLs embedded in shared tweets, in this study, we only include tweets that contain at least one URL. Detailed statistics on the number of tweets with URLs, the number of engaged users, and the distribution of tweet categories are provided in Table 1.
Identifying “Friends”.
We further refine our dataset to isolate a specific subset of target users. In line with the methodology of Nikolov, Flammini, and Menczer (2021), we establish a minimum inclusion criterion: we study only users who shared at least 10 URLs of any kind during the overall observation period (12 months). This threshold ensures that we concentrate on users with sufficient data to reliably assess their susceptibility. Applying this criterion, our data narrows down to 698,808 users for the Election dataset and 1,047,566 users for the Covid dataset.
We define two users as “friends” if they engage in active reciprocal interactions. This means both users need to actively engage with the other via retweets, quotes, or replies. Although using the followers network could provide a more robust definition of friendship, such data is challenging to obtain using the Twitter API (Martha, Zhao, and Xu 2013). In lieu of the followers data, we use interactions (retweet, quote, and reply) as a proxy for friendship, consistent with previous studies (Jiang, Ren, and Ferrara 2023; Rao, Morstatter, and Lerman 2022). After filtering for users in reciprocal friendships, we have 49,794 and 73,960 users from the Election and Covid datasets, respectively.
Constructing Friendship Networks.
We construct three distinct friendship networks based on the types of reciprocal interactions: Interaction, Retweet, and Mention networks. The Interaction network includes all forms of interactions between users, such as retweets, quotes, and replies. In the Retweet network, connections capture retweets only, whereas the Mention network includes interactions through replies or quotes. We differentiate the Retweet and Mention networks due to their varying social implications and levels of endorsement, a distinction further explored in the “Adoption” section. In Table 1, we present the number of nodes and edges for each network in the two datasets.
Results
RQ1: Susceptibility and Homophily
Figure 2: The susceptibility metrics (IAR and SAR) exhibit homophily within the social network, as evidenced by the weighted average correlation (Wgt. Avg. Corr.) between users and their friends. All correlation coefficients are significant ($p<0.001$).
Figure 3: Paradox holding probability $p(k,s)$ at varying degrees $k$ using the Interaction network in the Covid dataset.
We first examine whether susceptibility to influence exhibits homophily in a friendship network. To this end, we see if a user’s susceptibility metrics are related to their friends. Following Jiang et al. (2023), we measure homophily by calculating the Pearson correlation between each user’s susceptibility score and the weighted average susceptibility scores of all their friends. This method enables us to consider the frequency of user interactions and accounts for cases where users may have varying levels of influence across different connections. A coefficient of $-1$ suggests that users preferentially connect with others who have opposing susceptibility scores, while a coefficient of $1$ indicates a preference for connections between users with similar susceptibility scores.
The results presented in Figure 2 show that susceptibility is homophilous with respect to both the IAR and SAR scores, as indicated by the statistically significant and positive weighted average correlations. The results are consistent across all types of networks in both datasets. To further confirm the reliability of these findings, we compared our results with those from two null models. The first null model maintains the original network’s degree distribution by randomly selecting and swapping the endpoints of edge pairs. The second null model preserves only the number of edges by randomly reassigning neighbors to nodes at one end of each edge. Both null models result in weighted average correlation measures that are not significant, therefore confirming that our results on the real data cannot be attributed to statistical flukes. Please refer to the Appendix for details.
Our analysis reveals that there is indeed homophily in user susceptibility. Moreover, this homophily, which refers to the tendency for individuals to form connections with others who are similar to them, appears to be more prevalent in influence-driven adoption (IAR) than in spontaneous adoption (SAR). This suggests that within social networks, individuals connected by social ties tend to exhibit similar inclinations toward adopting the behaviors of their friends. This is likely due to the stronger impact of peer influence within close social circles, where interactions are more frequent and intense (De Klepper et al. 2010). As a result, a reinforcing loop emerges: individuals in tight-knit groups are more likely to adopt similar behaviors or attitudes due to the influence of their peers.
Summary.
Susceptibility to influence is homophilous. In particular, influence-driven behavior (as measured by IAR) appears to be more homophilous than spontaneous-driven activity (as measured by SAR).
RQ2: The Generalized Friendship Paradox of Susceptibility to Influence
Table 2: Empirical results for the Generalized Friendship Paradox in three social networks derived from the Election dataset are presented. For each susceptibility metric $s$ (IAR and SAR), we calculate the Spearman correlation coefficient $\rho_{ks}$ between degree $k$ and $s$ ($p<0.001$ for all $\rho_{ks}$), average paradox holding probability $P$ and the average susceptibility $\langle s\rangle$ of nodes and their friends $\langle s\rangle_{nn}$.
Network Metric $s$ $\rho_{ks}$ Individual-level GFP Network-level GFP
Real-Network
$P$
Baseline 1
$P$
Baseline 2
$P$
$\langle s\rangle$
Real-Network
$\langle s\rangle_{nn}$
Baseline 1
$\langle s\rangle_{nn}$
Baseline 2
$\langle s\rangle_{nn}$
Interaction IAR 0.179 0.541 0.672 0.798 0.007 $<$ 0.008 0.019 0.028
SAR --0.166 0.409 0.326 0.249 0.880 $>$ 0.868 0.810 0.751
Retweet IAR 0.298 0.501 0.662 0.824 0.013 $<$ 0.014 0.022 0.040
SAR --0.294 0.425 0.355 0.250 0.821 $>$ 0.809 0.722 0.670
Mention IAR 0.111 0.544 0.623 0.697 0.005 $<$ 0.006 0.009 0.015
SAR --0.099 0.383 0.354 0.325 0.896 $>$ 0.880 0.852 0.834
When asking whether the Generalized Friendship Paradox (GFP) holds for susceptibility to influence, we aim to investigate whether individuals’ friends are generally more susceptible than they are. Consistent with Eom and Jo (2014), in this study, we formulate the GFP at both the individual and network levels. At the individual level, a node exhibits GFP if its susceptibility metric is lower than the average of its friends. At the network level, the GFP holds if the average susceptibility metric of all nodes is less than the average susceptibility metric of their respective friends.
Individual-level GFP.
The individual-level GFP holds for one user $i$ if the following condition is satisfied:
$s_{i}<\frac{\sum_{j\in\mathcal{N}_{i}}s_{j}}{k_{i}},$ (1)
where $s_{i}$ is the susceptibility metric of user $i$, $\mathcal{N}_{i}$ represents the set of $i$’s friends, and $k_{i}$ is the degree of $i$. Essentially, it calculates whether a user’s susceptibility is less than the average of their friends. Based on this formulation, we define the paradox holding probability $p(k,s)$ as the proportion of nodes with degree $k$ and susceptibility metric $s$ that satisfies the condition in Eq. (1).
Network-level GFP.
At the network level, the GFP examines the phenomenon across the entire network. It is determined by comparing the average susceptibility metric of every node $\langle s\rangle$ with the average susceptibility metric of all their friends $\langle s\rangle_{nn}$, computed as follows:
$\langle s\rangle_{nn}=\frac{\sum_{i=1}^{N}k_{i}s_{i}}{\sum_{i=1}^{N}k_{i}},$ (2)
where a node $i$ with degree $k_{i}$ is counted as a neighbor $k_{i}$ times. If the average susceptibility metric of a node’s friends exceeds the node’s own average susceptibility metric ($\langle s\rangle<\langle s\rangle_{nn}$), then the GFP holds at the network level.
GFP: Comparing SAR vs. IAR.
Figure 3 depicts the paradox holding probability $p(k,s)$ as a function of degree $k$ and susceptibility metric $s$ using the Interaction network in the Covid dataset. Similar findings from the Election dataset are presented in the Appendix. This three-dimensional representation reveals that, with a constant degree $k$, $p(k,s)$ decreases as $s$ (IAR or SAR) increases, aligning with findings from Eom and Jo (2014). This implies that users with very low susceptibility $s$ tend to connect with those having higher $s$ values, resulting in a $p(k,s)$ close to $1$. Conversely, users with very high susceptibility $s$ values exhibit a $p(k,s)$ approaching 00. Users across this spectrum exhibit a consistent trend, wherein $p(k,s)$ remains close to 1 when $s$ and $k$ are strongly, positively correlated. This effect is particularly evident for IAR, indicating a specific pattern where the paradox holds.
Table 2 presents the empirical results of the individual and network-level GFP across three friendship networks constructed from the Election dataset. We get similar results from the Covid dataset. For each susceptibility metric $s$ (IAR or SAR), we calculate the average paradox holding probability on the individual level $P$, integrating $p(k,s)$ over $k$ and $s$, following Eom and Jo (2014), and the Spearman correlation coefficient between node degree $k$ and susceptibility metric $s$, denoted as $\rho_{ks}$. For the network-level GFP, we compare the node average $\langle s\rangle$ and their friends’ average $\langle s\rangle_{nn}$. We also employ two baseline models for comparison based on the null models described above.
For IAR, the value of $P$ exceeds 0.5 in every friendship network, indicating that the GFP holds at the individual level. Furthermore, the condition $\langle s\rangle<\langle s\rangle_{nn}$ is consistently satisfied across networks, indicating that the GFP also holds at the network level. This implies that, on average, individuals’ friends are more likely to adopt content seen on social media than the individuals themselves. In contrast, for SAR, $P$ consistently falls below 0.5, and $\langle s\rangle$ is always higher than $\langle s\rangle_{nn}$. These observations suggest that the GFP does not hold for spontaneous adoption. In simpler terms, individuals are less inclined to adopt unseen content compared to their friends.
Moreover, it is worth noting that while the GFP is observed for IAR, it is less pronounced compared to the two null models. The discrepancy between individual IAR and that of their friends is smaller in the observed network than in a randomly constructed network. For example, at the individual-level, in the Interaction network of the Election dataset, a random baseline model predicts that in 79.8% of cases, an individual’s friends, on average, would exhibit higher IAR than the individual. However, the actual observed rate in the real-world network is only 54.1%. At the network-level, the differences between $\langle s\rangle$ and $\langle s\rangle_{nn}$ are also smaller in the actual network. This reduction in the discrepancy between a user’s susceptibility and their friends’ average susceptibility is similarly observed for SAR.
These observations corroborate the study by Lerman, Yan, and Wu (2016), which demonstrates that the “majority illusion,” a phenomenon resulting from the friendship paradox, is more intense within networks characterized by heterogeneous degree distributions and disassortative structures. Our observations of real-world social media networks, which exhibit greater homophily compared to random networks, indeed show a reduced paradox holding probability. This implies that the structural attributes of a network can significantly impact the manifestation of the GFP.
Summary.
Our findings indicate that the Generalized Friendship Paradox holds for influence-driven behavior but not for spontaneous action in friendship networks. This suggests that users’ friends tend to be more susceptible to influence and less spontaneous than the users themselves. Interestingly, the paradox holding probability and the differences in IAR and SAR between users and their friends are less noticeable compared to the random baselines; the less homophilous the network, the larger the discrepancy.
RQ3: Predicting Susceptibility to Online Influence
As the GFP provides valuable insights into the relation between an individual’s susceptibility and that of their friends, and both IAR and SAR are homophilous, we investigate the potential to predict a user’s susceptibility based on their friends’ susceptibility levels. Furthermore, we evaluate whether incorporating additional features, such as user and friends’ metadata along with network properties, can enhance prediction accuracy. To accomplish this, we employ both simple linear regression and a random forest model.
Linear Regression Model.
First, we assess the effectiveness of predicting a user’s IAR (resp. SAR) based on the average IAR (resp. SAR) of their friends. The linear regression model is expressed as:
$s_{user}=\beta_{0}+\beta_{1}\times\bar{s}_{friend}+\epsilon,$ (3)
where $s_{user}$ and $\bar{s}_{friend}$ refer to a user IAR (resp. SAR) and their friends’ average IAR (resp. SAR).
Table 3 shows the results of the simple linear regression analysis. From Table 3, we note that both IAR and SAR models yield positive regression coefficients ($\beta_{1}$), suggesting a positive relationship between a user’s susceptibility metrics and those of their friends. Notably, the $R^{2}$ values for the IAR models are consistently higher than those of the SAR models, indicating that influence-driven adoptions are more accurately explained by friends’ susceptibility compared to spontaneous adoptions.
Additionally, we also attempt to predict user IAR using friends’ SAR and vice versa; however, these models produce $R^{2}$ values all below 0.05 in both datasets, highlighting that although strongly negative-correlated, IAR and SAR do not adequately predict each other across social networks.
Table 3: Summary of linear regression results for predicting user susceptibility metrics based on friends’ average susceptibility metrics. All regression coefficients reported are significant ($p<0.001$).
Dataset $s_{user}$ $\bar{s}_{friend}$ $\beta_{0}$ $\beta_{1}$ $R^{2}$
Election IAR IAR 0.002 0.622 0.363
Election SAR SAR 0.550 0.377 0.137
Covid IAR IAR 0.005 0.533 0.283
Covid SAR SAR 0.537 0.392 0.134
Random Forest Model.
To assess whether a more advanced model equipped with a broader range of features can enhance predictive accuracy, we employ an interpretable ensembling machine learning algorithm. This approach also aims to enhance our understanding of the factors that impact users’ susceptibility to social influence. Specifically, we utilize a random forest regression model that incorporates not only susceptibility metrics but also network properties and user metadata. This features set includes friends count, followers count, favorites count (the total number of tweets a user has liked), and statuses count (the total number of tweets a user has posted), along with network properties from the Interaction network such as degree centrality, eigenvector centrality, and clustering coefficient.
We divide users into training and testing sets and use the variance inflation factor (VIF) to assess multicollinearity among the features. In both the Election and Covid datasets, all VIF values are below 5, indicating no significant multicollinearity issues (James et al. 2013). We optimize the random forest regressor using a randomized search on the hyperparameter space to efficiently predict user IAR and SAR scores. We configure the randomized search to sample 100 different parameter settings using 5-fold cross-validation. After identifying the best parameters (see the Appendix for more details), we evaluate each model on a test set using $R^{2}$.
Table 4 provides an overview of the results. The random forest regressor shows promising performance, particularly for the IAR metric. It slightly outperforms the linear regressor, as indicated by the $R^{2}$ difference labeled as $\Delta R^{2}$ in Table 4. This difference is more pronounced for the SAR metric, while it remains quite small (less than 0.05) for the IAR metric, suggesting that IAR can be effectively predicted using solely friends’ IAR, as previously demonstrated in the linear regression model. A study on feature importance further supports this intuition.
Table 4: Summary of the performance of the random forest and linear regressors. $R^{2}$ denotes the coefficient of determination.
Election Covid
Susceptibility Metric IAR SAR IAR SAR
Random Forest $R^{2}$ 0.406 0.389 0.332 0.312
Linear Regression $R^{2}$ 0.363 0.137 0.283 0.134
$\Delta R^{2}$ 0.043 0.252 0.049 0.178
Figure 4: Top 5 random forest feature importances in the Covia dataset for IAR and SAR prediction.
Figure 4 illustrates the feature importance (top 5 features) for the random forest regressor in predicting IAR and SAR within the Covid dataset. Notably, friends’ IAR emerges as the most significant feature for predicting IAR, with other features proving substantially less relevant. This finding confirms our previous intuition and remains consistent in the Election dataset. For SAR predictions, while friends’ SAR is the most critical feature, additional factors such as statuses count, friends’ IAR, and favorites count also contribute. This pattern holds in the Election dataset, where statuses count is the most important feature.
Figure 5: SHAP value distributions for SAR predictions in the Covid dataset. Each distribution illustrates the direction and magnitude of the impact of individual features on the model’s output, with features ordered by importance from top to bottom.
Output Explainability.
To further elucidate the results from feature importance and quantify the specific contribution of each feature to individual predictions, we compute the SHAP (SHapley Additive exPlanations) values (Lundberg and Lee 2017) of the dependent variables predicted in the test set. SHAP values provide a nuanced measure of the impact of each feature on individual instances. Each feature is assigned a SHAP value per instance, which can be aggregated to offer a comprehensive view of the direction and magnitude of its impact on the model’s predictions. Figure 8 shows a SHAP summary plot for the SAR metric in the Covid dataset (additional results are detailed in the Appendix). On the y-axis, features are listed in order of impact, while the x-axis displays the SHAP values, indicating the change in log odds caused by each feature. Points represent individual data instances, colored red (resp. blue) for high (resp. low) feature values.
The most impactful features affecting the SAR predictions in the Covid dataset are friends’ SAR, user statuses count, and user clustering coefficient. The features friends’ SAR and user statuses count have significant opposing impacts on the predictions, indicating that these features can notably increase or decrease the SAR value. Specifically, high values of friends’ SAR tend to increase the SAR score, while higher values of user statuses count tend to decrease it. This relationship between user statuses count and SAR suggests that a high volume of shared tweets correlates with a small rate of spontaneous adoptions. Interestingly, the user clustering coefficient also emerges as a significant factor, where higher clustering coefficients typically decrease SAR, possibly indicating that users with denser social networks may experience fewer spontaneous adoptions. This finding is consistent in the Election dataset (see the Appendix), where user statuses count, friends’ SAR, and user favorites count are the most impactful features.
For IAR prediction (see the Appendix), the most impactful feature is friends’ IAR, with high values of friends’ IAR significantly increasing the IAR score, supporting our previous results related to homophily in IAR.
Summary.
Users’ level of influence-driven adoption can be effectively predicted by solely leveraging friends’ IAR. In contrast, the predictability of a user’s SAR is significantly enhanced by incorporating additional features, such as counts of tweets and likes. Explainable models reveal that SAR is affected by a wider array of factors, suggesting a more complex interplay of user engagement and preferences in the spontaneous adoption of content.
Discussion and Conclusions
This study, based on the analysis of millions of tweets, explores the networked dynamics of user susceptibility to online influence from the dual perspective of influence-driven and spontaneous adoption. We present four key insights from our research.
Finding 1: Susceptibility is homophilous.
We find that susceptibility is homophilous in social networks, as evidenced by the positive correlations between users’ adoption rates (either influence-driven or spontaneous) and those of their friends. This indicates that individuals prone to influence often connect with similarly susceptible peers. Spontaneous adoption shows significant but lower homophily.
Finding 2a: The Susceptibility Paradox holds for influence-driven behavior.
We observe the Generalized Friendship Paradox (GFP) for user susceptibility to influence, termed the Susceptibility Paradox, across two dimensions: First, individuals in a social network tend to be connected to others who are more likely to adopt content due to social influence, as reflected in higher IAR; second, they engage in less spontaneous content adoption, as indicated by lower SAR. Collectively, these observations suggest that “your friends are generally more susceptible to social influence than you.” Additionally, we observe that IAR and SAR are negatively correlated, suggesting that users with higher influence-driven adoption are less likely to share content spontaneously.
Finding 2b: The Susceptibility Paradox is less pronounced when homophily is present.
While we observe that the GFP holds for IAR but not for SAR, both IAR and SAR demonstrate homophily. These homophilous patterns indicate that individuals are prone to connect with others who exhibit similar rates of influence-driven and spontaneous adoption. This presents two countervailing forces where, although one’s friends typically have higher IAR but lower SAR than oneself, they tend to remain similar in terms of overall susceptibility scores. Comparing the individual- and network-level Generalized Friendship Paradox in homophilous real-world networks versus non-homophilous baseline networks, our findings suggest that the GFP is less pronounced in homophilous networks, where high-IAR (or SAR) nodes tend to connect with other high-IAR (or SAR) nodes. This corroborates Lerman, Yan, and Wu (2016)’s finding that the “majority illusion” is more pronounced in heterogeneous, disassortative networks.
Finding 3: IAR can be predicted by friends’ IAR and is more predictable than SAR.
Our analysis indicates that IAR can be accurately predicted solely based on friends’ IAR, whereas SAR is influenced by a broader range of factors. This distinction adds depth to our understanding of susceptibility to influence. On the one hand, our results suggest that influence-driven behaviors (IAR) tend to follow more predictable patterns. This likely stems from the strong link between influence-driven adoption and observable network structures, such as homophily (RQ1) and GFP (RQ2). For instance, studies have shown that targeted marketing and influence campaigns can be effectively strategized based on previous adoption patterns of a consumer’s network (Hill, Provost, and Volinsky 2006). On the other hand, spontaneous behaviors (SAR) are influenced by a myriad of internal and external, often stochastic factors, making them less predictable.
Implications.
Our research offers practical insights for social media platforms, policymakers, and marketers. The observed homophily in susceptibility suggests that targeted interventions in fields like public health and marketing can effectively target influenceable audiences. By identifying these cohesive, vulnerable groups, intervention strategies can be tailored to mitigate the spread of harmful content and encourage positive behaviors more efficiently (Centola 2011). Further, the observation that the Generalized Friendship Paradox applies to influence-driven adoption could inspire efficient sampling algorithms to identify highly susceptible populations (Alipourfard et al. 2020). Given that IAR is more predictable and that individuals with high IAR are more influenced by their social connections, platforms can specifically target vulnerable audiences by nudging them with accurate, fact-checked content to prevent the endorsement and spread of misinformation.
Limitations.
We acknowledge several limitations in our research. First, our study draws on datasets specifically related to the 2020 US election and the COVID-19 pandemic, capturing only a subset of users over an extended but limited period. Consequently, generalizing our results to broader user networks, different contexts, or other social media platforms necessitates further validation. Second, we infer exposure based on observable user interactions, which might not comprehensively capture exposures within Twitter and across different media. Nonetheless, this method for assessing exposure was tested and validated in prior research (Ferrara and Yang 2015; Rao, Morstatter, and Lerman 2023; Ye et al. 2024). Third, although friends’ IAR explains 36% and 28% of the variation of user IAR in Election and Covid datasets, respectively, we recognize that social influence is a complex phenomenon and may be influenced by various unobserved personal, societal, and cultural factors. Despite these limitations, the consistency of our results across multiple intertwined analyses on distinct datasets supports the reliability and robustness of our findings.
Ethical Statement.
Throughout our research process, we have adhered to stringent ethical standards. We have ensured that no personally identifiable information was utilized, and all analyses were conducted on aggregated data. Additionally, we have thoroughly considered the societal impacts of our research. We acknowledge the potential misuse of our findings by malicious entities to target vulnerable individuals susceptible to manipulation. However, we firmly believe that the significance of our work outweighs these risks. Our research has the potential to inform the development of online systems that prioritize safety and responsibility for all individuals, thereby contributing positively to the integrity of digital environments.
Appendix
Frequency Distributions of IAR and SAR
Figure 6 shows the frequency distributions of the IAR and SAR metrics in Election and Covid datasets.
Figure 6: Frequency distributions of the IAR and SAR metrics in Election and Covid datasets.
Explanation of the Null Models
Baseline 1: Edge Swapping
This null model is designed to test network robustness while preserving the original network’s degree distribution and total number of edges. The model operates by implementing an edge swapping mechanism: (1) Pairs of edges are randomly selected within the network. Care is taken to ensure that these selections do not result in multiple edges between the same pair of nodes (unless the original network allows for multiple edges) or self-loops; (2) The endpoints of these selected edge pairs are swapped—this means that if edges $(u,v)$ and $(x,y)$ are chosen, they might be swapped to $(u,y)$ and $(x,v)$; (3) This process is repeated multiple times to ensure thorough randomization while keeping the number of connections (degree) each node has intact.
Baseline 2: Random Neighbor Assignment
This model tests the robustness of network properties against random reassignment of edges: (1) Each node in the network retains its degree but has its neighbors randomly reassigned. This means that while each node will still have the same number of connections, the actual connections (neighbors) will be randomly chosen; (2) The total number of edges and the overall degrees are preserved, similar to Baseline 1. However, unlike edge swapping, this method completely disrupts any existing neighbor relationships. The objective here is to create a completely randomized structure where only the degree sequence of the original network is preserved.
Paradox holding probability $p(k,s)$ at varying degrees $k$
Figure 7 depicts the paradox holding probability $p(k,s)$ as a function of degree $k$ and susceptibility metric $s$ using the Interaction network in the Election dataset.
Figure 7: Paradox holding probability $p(k,s)$ at varying degrees $k$ using the Interaction network in the Election dataset.
Random Forest Model Parameters
Parameters tested included the number of trees (“n_estimators”; 100 to 1000 in increments of 50), features per split (“max_features”; scaled linearly from 1 to the total number of features), maximum tree depth (“max_depth”; 10 to 100 and unlimited), and minimum samples for splits and leaves (“min_samples_split” and “min_samples_leaf”), using values of 2, 5, 10, and 1, 2, 4, respectively. Table 5 shows a summary of the random forest model parameters.
Table 5: Summary of the random forest model parameters.
Election Covid
Susceptibility Metric IAR SAR IAR SAR
n_estimators 750 800 750 800
max_features 4 4 4 4
max_depth 70 80 70 80
min_samples_split 5 10 5 10
min_samples_leaf 2 2 2 2
SHAP Values of IAR and SAR Prediction
Figure 8: SHAP value distributions for SAR predictions in the Election dataset. The plot illustrates the direction and magnitude of impact of individual features on the model’s output, with features ordered by importance from top to bottom.
Figure 8 shows the SHAP value distributions for SAR predictions in the Election dataset. Figure 9 show the SHAP value distributions for IAR predictions in the Election dataset.
Figure 9: SHAP value distributions for IAR predictions in the Election dataset. The plot illustrates the direction and magnitude of impact of individual features on the model’s output, with features ordered by importance from top to bottom.